Assume that an investor has a long position in standard CME EUR futures contract

(EUR 125,000) and that today’s settlement price is $1.1678 per euro. If the balance in

the investor’s performance bond account is currently $2,500 and according to his/her

broker’s margin rules new money must be deposited if the balance falls to $1,900, at

what EUR futures settlement prices would his/her broker demand additional money

be deposited?
Answer #1

Initial account balance is $2,500 and investor has a long position on 125,000 Euro at a price of $1.1678.

We need to find the closing price (CP) at which a loss of $600 will be incurred, so the balance of the account is $2,500 - $600 = $1,900.

($1.1678 – CP) (125,000) = $600.

CP = 1.1678 - (600/125000) = 1.163

Therefore, if CP falls below 1.163, the investor will receive a margin call, and will need to deposit additional money